### v2

# The Tautochrone of Huygens and Abel: From Constructive Geometry to Fractional Calculus

Abstract
In this paper, we explore the connections between Christiaan Huygens and Niels Henrik Abel through the tautochrone problem. The problem — determining the curve along which a particle descends under gravity in the same time, regardless of its starting point — has been a central topic at the intersection of physics, geometry, and analysis. Though these two major figures are separated by nearly two centuries, they approached the problem in radically different ways. While Huygens proposed a physical solution based on geometric construction, Abel approached the problem within the analytic framework of integral equations, employing a procedure that can be seen as anticipating and paving the way for the development of differential calculus of arbitrary order. This contrast highlights a broader historical narrative: the transformation of mathematical thinking from constructive geometry to abstract analysis.
Keywords: Isoperimetric problems; variational calculus; integral equations; fractional calculus
I Introduction
Problems involving extremes (maximums and minimums) are old and persistent. The very use of terms like “law of least effort”, “path of least resistance”, “shortest distance between two points”, etc., is part of our common sense of economy or simplicity in the actions of nature Fermat . The first problem of finding an extreme that is known to us is associated with the name of the Phoenician Princess Dido, queen of Carthage — immortalized by the Latin poet Virgil in the Aeneid, and later by the English composer Henry Purcell (1659–1695) in the musical drama Dido & Aeneas. The legend says that the princess, having emigrated to North Africa after being defrauded of her belongings by her brother, Pygmalion, was promised by Hiarbas, the native chief of the region, as much land as she could enclose in a cowhide. Cleverly, she cut the cowhide into a very long strip and placed it between herself and the sea, creating a territory large enough to build the city of Carthage Kelvin . The typical solution found by Dido is to arrange the strip in a circular shape. It is, therefore, an isoperimetric problem, one of many others that will be encountered throughout history. Another problem of this nature is associated with the name of Heron of Alexandria (10–70 A.D.). He is credited with deriving the law of reflection from the principle that the light ray emitted from point A reaches point B after being reflected by a mirror, following the path that takes the least time. This principle was extended by Fermat to derive Snell–Descartes’ law of refraction Fermat .
In this work, we focus on a problem of this nature, which also belongs to the realm of isoperimetric and optimization problems — the tautochrone problem. This problem concerns the determination of the curve along which a particle, starting from any point on the curve under the gravitational force, will reach the lowest point in the same amount of time, regardless of its initial position. The tautochrone problem is closely related to the concept of extremal times and optimal paths, much like the problems that Dido and Heron of Alexandria faced. In fact, solving the tautochrone problem is akin to finding an optimal path that minimizes a certain quantity — in this case, the time of descent — given constraints. It’s another example of how nature often seeks optimization, whether through the geometry of curves or the behavior of physical systems under certain conditions.
The problem was also posed by the mathematician Christiaan Huygens in the 17th century and was later shown to have a solution in the form of the cycloid. What Huygens demonstrated was that a material point, released from rest and allowed to slide without friction along an arc of an inverted cycloid, reaches the lowest point of the curve in the same amount of time, regardless of the initial height. Thus, a pendulum constrained to oscillate along such a cycloidal arc exhibits a period of oscillation that is truly independent of the amplitude of its motion. Huygens’ approach was fundamentally constructive. He did not yet have access to the fully developed calculus of Newton and Leibniz, but relied instead on geometric arguments and proportions, integrating physical intuition with mathematical ingenuity. His solution was elegant, self-contained, and tailored to a particular problem with practical consequences.
By contrast, in the early 19th century, Niels Henrik Abel approached a more general class of problems: inverse integral equations. In a brief but brilliant life, Abel transformed how mathematicians viewed such problems. He considered equations of the form
$\psi(a)=\int_{x=0}^{x=a}\frac{f(x)}{\sqrt{a-x}}\mskip 3.0mudx,$ (1)
and posed the question: given $\psi(a)$, can one determine the function $f(x)$? This type of integral, now known as an Abel integral, has direct analogies to the time-of-descent problem in the tautochrone curve, where the time taken to reach a point is an integral over a function of the height difference. Abel’s treatment of this inverse problem marked a turning point. Instead of seeking a specific curve, he found a general method for recovering unknown functions from their integrals — paving the way for later developments in functional analysis and integral transforms. The tautochrone problem became, in his hands, a special case of a much broader theory.
Here, we aim to highlight the scientific contributions of these two mathematicians, using the tautochrone problem as a guiding thread. Some biographical details are included to underscore the contrast between their backgrounds, beginning in childhood, when both had already shown a rare and precocious talent for mathematical science and abstract thinking. By focusing on how each of them approached the problem — separated by two centuries — we also trace how analysis became increasingly sophisticated in addressing important and advanced problems in physics and mathematics. Special attention is given, in the final part of the paper, to the manners in which Abel’s work paved the way for the development of key concepts in fractional calculus — matured in the years that followed Ross1977 .
II The Geometric Construction of Huygens
Christiaan Huygens was born on April 14, 1629, in Den Haag (The Hague), in the Dutch Republic. He was the son of Constantijn Huygens, a diplomat with a strong background in philosophy and the natural sciences. Christiaan’s father was also a distinguished poet, securing an enduring place in the history of Dutch literature. Constantijn ensured that his son received an exceptional education and had access to the leading intellectual circles of the time. A prominent role in this upbringing is attributed to Father Marin Mersenne (1588–1648) — who corresponded regularly with Huygens’s father — as well as to the friendship the elder Huygens maintained with René Descartes (1596–1650). Educated at home by private tutors until the age of sixteen, Christiaan benefited from the intellectual environment surrounding his family. Descartes, who was then living in the Dutch Republic and maintained a friendly relationship with Constantijn, is believed to have taken interest in the young man’s mathematical development. In 1645, Huygens enrolled at the University of Leiden, where he studied law, mathematics, and classical languages. Two years later, in 1647, he continued his legal studies at the College of Orange in Breda. The period from 1650 to 1666 would prove to be the most fertile and productive phase of his scientific life.
II.1 The Tautochronism of the Cycloid
Because astronomical investigations require increasingly accurate time measurements, Huygens devoted considerable effort to the problem. In 1656, he patented the first pendulum clock, significantly improving the precision of timekeeping. This innovation was later described in his 1658 publication Horologium (“The Clock”). Although his investigations into centrifugal force also date back to this period (1659), it was in his major work of 1673, Horologium Oscillatorium sive de motu pendulorum (“The Pendulum Clock, or On the Motion of Pendulums”), that he developed a full theoretical account of pendular motion Huygens1673 . It is also in this treatise that the modern formula for the centrifugal force in uniform circular motion first appears. The Horologium Oscillatorium is often regarded as one of the most important work on mechanics prior to Newton’s Principia (1687). In the book, Huygens derived the formula for the period of a simple pendulum under the small-angle approximation — a classical result encountered in the Galileo’s work. Huygens went further by investigating the precise curve along which a pendulum would swing with a period that is truly independent of its amplitude. His solution identified this path as a cycloid. In a few words, we can say that Huygens showed the cycloid is a tautochrone in a uniform gravitational field.
Figure 1: This is the original figure presented in the Huygens work in connection with the results analyzed in Propositions XXV–XXVI. The particle is released from the point B and the final point of its trajectory is in A. We notice that EA is parallel to BG and both describe a kind of inclined plane for the descend of the particle.
1
Indeed, this remarkable property is known as the tautochronism of the cycloid — from the Greek words ταυτό (equal, same) and ϰρόνος (time) — and it requires that the pendulum follow an inverted cycloidal trajectory. The tautochrone problem is addressed in Part II, specifically in Propositions XVI through XXVI. To reproduce his construction, we need the help of Fig. 1, part of the Huygens’ book Huygens1673b . For a better understanding of the discussion that follows, here is a summary of the elements that appear in Fig. 1 and their meaning:
  • Point A — is the lower vertex of the cycloid, the endpoint of the motion;
  • BA is an arc of the cycloid — the path followed by a body in free fall;
  • Vertical axis DA — a straight line along which the ‘standard’ time is measured for comparison of the motion;
  • Semicircle FHA — constructed on the diameter FA, used to establish the proportion of times;
  • Tangent BG — the line tangent to the cycloid at point B, parallel to the line EA, which intersects the semicircle.
Let us now proceed by sketching some of the representative steps of his derivation, focusing the last three propositions, starting with a brief summary of the Proposition XXIV, from which follow almost immediately the Propositions XXV and XXVI. This is useful because Huygens’ solution rests on a specific geometric property of the cycloid, established in Proposition XXIV, which he then applies in the culminating Propositions XXV and XXVI. In this perspective, Proposition XXIV is the geometric foundation of his derivation because, in this framework, he first proves a critical lemma: for any point on an inverted cycloid, the ratio of the arc length from that point to the vertex to the straight chord connecting them is constant. More precisely, for any point I on the cycloid arc BA (see Fig. 1),
$\frac{\text{arc}IA}{\text{chord }IA}=\frac{\text{semicircumference of the generating circle}}{\text{its diameter}}=\frac{\pi}{2}.$
This constant arc–to–chord ratio is a kind of kinematic signature of the cycloid. It is not an obvious identity; Huygens derives it through an intricate geometric construction involving the cycloid’s evolute Huygens1673b . This proposition provides the precise geometric tool needed to translate distances into times.
Using the geometric identity established in Proposition XXIV, Huygens proves the tautochrone property in Proposition XXV:
Proposition XXV In a cycloid with a vertical axis, and with the vertex seen to be the lowest point, the times of descent for some body, on leaving any point on the cycloid from rest until it reaches the lowest point at the vertex, are equal to each other; and this time has the same ratio to the time of fall along the whole axis of the cycloid as the semi-circumference of the circle to the diameter.
In the accompanying construction (Fig. 1), Huygens demonstrates that the descent time along the cycloid arc BA is proportional to the arc length of the auxiliary semicircle FHA, while the reference free-fall time along the axis DA is proportional to its diameter FA Bruce . The ratio between these times—the constant $\pi/2$ from Proposition XXIV—thus quantifies the tautochrone period: it equals $(\pi/2)$ multiplied by the free-fall time along the axis.
The kinematic reasoning proceeds as follows: the time along arc BA relates to the time of uniform motion along line BG (with half the speed acquired in free fall) as arc FHA relates to line FA. By Galileo’s laws Galileo , that uniform motion time equals the free-fall time along DA. Consequently, the cycloidal descent time is permanently tied to the $\pi/2$ ratio of the generating circle.
Huygens then examines the internal distribution of this constant time in Proposition XXVI Bruce :
Proposition XXVI With the same positions, if some line HI is drawn above which cuts the arc BA in I, and the circumference FHA in H: I say that the time to pass through the arc BI, to the time to cross the arc IA after BI, has the ratio which the arc of the circumference FH has to HA.
This proposition reveals the geometric clockwork inside the isochronous motion: time elapsed maps linearly onto the arc of the generating circle. For any intermediate point I, the ratio of times (BI to IA) equals the ratio of circular arcs (FH to HA). In modern terms, this expresses the differential condition that ensures the total time remains constant; each segment of the descent contributes a duration proportional to a corresponding arc segment on the auxiliary circle.
Thus, Huygens’s synthetic method follows a clear architecture:
  • Proposition XXIV provides the fundamental geometric identity (arc/chord = $\pi/2$).
  • Proposition XXV uses this identity to prove and quantify global isochrony.
  • Proposition XXVI derives the internal temporal structure, confirming the self-consistency of the geometric clock.
Crucially, Huygens did not derive the cycloid from the tautochrone condition. Instead, starting from the known curve (generated by a rolling circle of radius $r$ and then inverted), he verified its isochrony through geometric proportions and kinematic equivalences. His reasoning was strictly comparative—relying on Galileo’s laws and the fixed ratio
$\frac{\text{arc IA}}{\text{chord IA}}=\frac{\pi}{2},$
and never invoked differential equations or variational principles. What a modern reader might interpret as the curve’s curvature adjusting tangential acceleration is, for Huygens, the consequence of a geometric proportion between arcs, chords, and times.
II.2 From Huygens’s Geometry to Modern Analytic Language
Huygens’s 1673 demonstration of the tautochrone property was purely geometric. As we have seen, he established that for an inverted cycloid, the descent time from any point to the vertex equals the constant $T=\pi\sqrt{r/g}$, where $r$ is the radius of the generating circle. His proof relied entirely on proportions between arc lengths, chords, and times—never on analytic expressions.
At the heart of Huygens’s reasoning is a kinematic correspondence: the descent along the cycloid arc BA maps linearly onto uniform motion along the arc of its generating circle. Specifically, he shows that the vertical height of the particle on the cycloid corresponds to the angular position of a point moving uniformly along the auxiliary circle. Since uniform circular motion is isochronous, the descent along the cycloid inherits the same property. This geometric insight also explains why constraining an ordinary pendulum with “cycloidal cheeks” renders its period amplitude-independent Huygens1673 ; Bruce .
In modern analytic terms, the tautochrone condition can be expressed as follows. For a particle sliding without friction from rest at height $y_{0}$, conservation of energy gives $v=\sqrt{2g(y_{0}-y)}$. The descent time is therefore
$T=\int_{0}^{y_{0}}\frac{ds}{\sqrt{2g(y_{0}-y)}}=\text{constant},$ (2)
independent of $y_{0}$. With $ds=\sqrt{1+(dy/dx)^{2}}\mskip 3.0mudx$, this becomes
$T=\int\frac{\sqrt{1+y^{\prime 2}}}{\sqrt{2g(y_{0}-y)}}\mskip 3.0mudx,$ (3)
taken along the unknown curve $y(x)$.
This formulation reveals the tautochrone as an integral constraint to be inverted—exactly the approach Abel would take in 1823 Abel1823 . For Huygens, however, the cycloid was not derived from this equation but verified against it through geometric means. His result can be summarized as
$\frac{T_{\text{cycloid}}}{T_{\text{free fall}}}=\frac{\pi}{2},$
a ratio that emerged from the geometry of the cycloid and its generating circle, not from manipulating integrals.
Connection to the Brachistochrone and Calculus of Variations.
A different but related problem—the brachistochrone (from Greek βρἑϰιςτος, shortest, and ϰηρῗνος, time)—was posed by Johann Bernoulli in 1696, two decades after Huygens’s work Bernoulli . Bernoulli asked for the curve of fastest descent between two given points (not equal-time descent from any starting point). The brachistochrone functional,
$T_{\text{brach}}=\int\frac{\sqrt{1+y^{\prime 2}}}{\sqrt{2gy}}\mskip 3.0mudx,$
mathematically resembles the tautochrone integral, though the problems are distinct: the brachistochrone minimizes time between two fixed points, while the tautochrone demands time invariance for variable starting points. Remarkably, both yield the same curve—the cycloid.
Jakob Bernoulli’s 1697 solution to the brachistochrone Bernoulli , followed by Euler’s systematization in 1744 Euler1744 ; Fraser1994 and Lagrange’s $\delta$-algorithm in the 1760s Lagrange-LF ; Fraser-LF , laid the foundations of the calculus of variations. This analytic development stands in stark contrast to Huygens’s geometric synthesis, highlighting the profound transformation of mechanics between 1673 and 1823—from synthetic verification to analytic derivation.
III The Integral Equations of Abel
Niels Henrik Abel was a Norwegian mathematician renowned for his groundbreaking work in algebra. He is best known for proving that the general equation of the fifth degree (quintic equations) cannot be solved by radicals, a result now known as Abel’s impossibility theorem (Abel–Ruffini theorem) Abel1824 ; Abel1826b . Abel’s work laid the foundation for much of modern algebra and group theory. Despite dying young at 26, Abel made significant contributions to mathematics, including studies on elliptic functions and the development of Abelian integrals. His legacy is celebrated in the mathematical community, and the prestigious Abel Prize is named in his honor.
Abel was born on August 5, 1802, in Nedstrand, Norway, into a modest family. His father, a Lutheran pastor, struggled with alcoholism, which led to financial difficulties. Despite this, Abel showed early signs of exceptional intelligence. He received his education at the local school in his hometown before moving to the University of Christiania (now Oslo) in 1821. Abel was initially enrolled in theology, as his family hoped he would follow in his father’s footsteps as a priest. However, he soon turned to mathematics, developing a deep interest in the subject, particularly in the areas of algebra and functions. His talent in mathematics became clear to his professors, and he quickly gained a reputation as a prodigy. Abel’s time at the university was financially strained, and he often had to struggle with limited resources. Despite this, he completed his studies, and by 1823, he sent his first important mathematical paper to the renowned mathematician Carl Friedrich Gauss. Gauss was impressed by Abel’s work, leading to a correspondence that would later have a significant influence on his career. Abel’s education was largely self-directed, supplemented by his interactions with prominent mathematicians of the time. However, due to financial hardship, he was unable to secure a formal academic position. As a result, he spent the last years of his life in poverty, moving between various cities in Europe, where he worked on his mathematical research. In 1825, Abel moved to Paris, where he met several important mathematicians, including Joseph Fourier, and continued his studies. Unfortunately, his health began to deteriorate, and he died of tuberculosis on April 6, 1829, at the age of 26. Despite his short life, Abel’s contributions to mathematics were profound and continue to influence the field to this day. His early life was marked by personal and financial struggles, but his mathematical brilliance shone through, leading to the development of key theories in algebra and analysis. This is completely different from the early life of Christiaan Huygens, who grew up in good health, in a financially comfortable and intellectually rich household, where the atmosphere was stable, cultured, and conducive to learning.
In 1823, Abel solved the tautochrone problem by using a fractional integration of order $\alpha=1/2$ Abel1823 and, in 1826, extended the approach to $\alpha\in(0,1)$ Abel1826 . In his 1826 work, Abel presents a different solution method, based on the use of the properties of Euler’s gamma function. As emphasized in Ref. Podlubny-Magin , it is a more concise and elegant method which, however, abandons the path opened toward fractional calculus, characteristic of his 1823 paper. Indeed, as we briefly discuss now, the solution proposed by him to the tautochrone problem involves the pioneering use of an integral operator of arbitrary order, and is the first explicit solution known of an integral equation Lutzen1982 .
III.1 Abel’s Solution: A Remarkable Theorem
Abel opens his 1823 paper with a statement that highlights his methodological shift Abel1823 :
It is well known that, with the help of definite integrals, many problems can be solved which otherwise cannot be resolved, or at least are very difficult to handle […] I will demonstrate a new application by solving the following problem.
He considers a generalization of the tautochrone problem, illustrated in Fig. 2. A particle descends from rest at point C along an unknown curve CA. Let $AB=a$ be the vertical height, and denote by $s(x)$ the arc length as a function of the horizontal coordinate $x=AP$. The descent time $T$ is prescribed as a given function $T=\psi(a)$. The problem is to determine the curve $s(x)$ that yields this time.
Figure 2: The illustrative picture appearing in the French translation of the Abel’s 1823 paper Abel1823 .
Abel derives the integral equation
$\psi(a)=\int_{x=0}^{a}\frac{ds}{\sqrt{a-x}},$ (4)
where $ds=s^{\prime}(x)\mskip 3.0mudx$. Instead of solving this specific case directly, he treats the more general equation
$\psi(a)=\int_{0}^{a}\frac{ds}{(a-x)^{n}},\qquad 0<n<1,$ (5)
with $\psi(a)$ finite at $a=0$.
After a series of manipulations involving power–series expansions and properties of the gamma function, Abel arrives at what he calls a “remarkable theorem” Podlubny-Magin :
$s(x)=\frac{\sin n\pi}{\pi}\mskip 3.0mux^{n}\int_{0}^{1}\frac{\psi(xt)\mskip 3.0mudt}{(1-t)^{1-n}}.$ (6)
In modern notation, this solution can be written as
$s(x)=\frac{1}{\Gamma(1-n)}\frac{d^{-n}}{dx^{-n}}\mskip 3.0mu\psi(x),$ (7)
which Abel recognizes as a fractional-order integral — an operation inverse to fractional differentiation. Indeed, he later writes:
If the equation of a curve is $s=\psi(x)$, the time a body takes to travel along an arc of the curve whose height is $a$ is $\sqrt{\pi}\mskip 3.0mu\dfrac{d^{1/2}\psi(a)}{da^{1/2}}$.
For the tautochrone ($n=1/2$), the solution reduces to
$\psi(x)=\sqrt{\pi}\mskip 3.0mu\frac{d^{1/2}}{dx^{1/2}}\mskip 3.0mus(x),$ (8)
which inverts the original integral. Crucially, Abel treats fractional integration and differentiation as mutually inverse procedures — a conceptual leap that foreshadows the later development of fractional calculus (see Refs. Liouville1832b ; Liouville1832a ; Liouville1834 ).
Equation (5), rewritten as
$\psi(t)=\int_{0}^{t}\frac{s^{\prime}(x)}{(t-x)^{n}}\mskip 3.0mudx,$ (9)
is — up to a constant factor — precisely what would later be called the Caputo fractional derivative of order $n$ Caputo1967 ; Caputo1969 . Abel’s solution therefore represents the first systematic inversion of such an integral equation, establishing a direct link between the tautochrone problem and the birth of fractional calculus.
III.2 Abel’s Isochrone: Explicit Solution
To derive Abel’s result in simpler terms, consider a particle of mass $m$ sliding from rest at height $y_{0}$ along a frictionless curve $s(y)$, as shown in Fig. 3. The descent time must be independent of $y_{0}$.
Figure 3: A mass $m$ under gravity slides down along a frictionless wire, whose figure is defined by $s(y)$, such that $s(0)=0$. This trajectory is equivalent to the arc BA in Fig. 1, rebuild here with more appropriate notation.
Conservation of energy gives $v=\sqrt{2g(y_{0}-y)}$, so that
$dt=-\frac{ds}{\sqrt{2g(y_{0}-y)}}=-\frac{1}{\sqrt{2g(y_{0}-y)}}\frac{ds}{dy}\mskip 3.0mudy.$
The total descent time is therefore
$T(y_{0})=\int_{0}^{y_{0}}\frac{1}{\sqrt{2g(y_{0}-y)}}\mskip 3.0mu\frac{ds}{dy}\mskip 3.0mudy.$ (10)
Equation (10) is an integral equation for $f(y)=ds/dy$. Writing $y_{0}\to y$ and $y\to z$, and noting that tautochrony requires $T(y)=\text{constant}/\sqrt{2g}$, we obtain
$k=\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$ (11)
where $k$ is constant. This is a convolution with kernel $K(y,z)=(y-z)^{-1/2}$.
The equation is readily solved via Laplace transforms. Using
$\mathcal{L}\left\{\frac{1}{\sqrt{y}};s\right\}=\sqrt{\frac{\pi}{s}},$
the transform of (11) yields
$\frac{k}{s}=F(s)\sqrt{\frac{\pi}{s}}\quad\Longrightarrow\quad F(s)=\frac{c}{s^{1/2}},$
with $c=k/\sqrt{\pi}$. Inverting gives
$f(y)=\frac{ds}{dy}=c\mskip 3.0muy^{-1/2}.$
Since $ds/dy=\sqrt{1+(dx/dy)^{2}}$, we have
$1+\left(\frac{dx}{dy}\right)^{2}=\frac{c^{2}}{y}.$
Setting $y=\frac{c^{2}}{2}(1-\cos\theta)$ leads to the parametric equations
$x$ $=\frac{c^{2}}{2}(\theta+\sin\theta),$ (12)
$y$ $=\frac{c^{2}}{2}(1-\cos\theta),$ (13)
which define a cycloid Sokolnikoff1958 . Thus Abel’s inversion of the integral equation recovers the same curve that Huygens had verified geometrically a century and a half earlier.
III.3 The Birth of Fractional Calculus
Let us pause for a moment to underline a central point of the present analysis. Abel’s solution to the tautochrone problem marks a pivotal moment in the history of mathematics: the emergence of fractional calculus. Equation (11),
$k=\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$
is not merely an integral equation; it defines a fractional integral of order $1/2$. Introducing the notation Davis :
$_{0}\mathrm{I}_{y}^{1/2}f(y)=\frac{1}{\Gamma(1/2)}\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$
the tautochrone condition becomes $_{0}\mathrm{I}_{y}^{1/2}f(y)=\text{constant}$. Solving it requires inverting this fractional operator — i.e., applying a fractional derivative $d^{1/2}/dy^{1/2}$.
Abel’s Insight: Inversion via Fractional Derivatives.
Abel recognized that if
$k={}_{0}\mathrm{I}_{y}^{1/2}f(y),$
then formally
$f(y)=\frac{d^{1/2}}{dy^{1/2}}\mskip 3.0muk.$
Thus, the problem reduces to computing the fractional derivative of a constant. In his 1826 paper Abel1826 , Abel solved the more general equation
$k(y)=\int_{c}^{y}\frac{f(z)}{(y-z)^{\alpha}}\mskip 3.0mudz,\qquad 0<\alpha<1,$ (14)
thereby introducing what would later be called the Riemann–Liouville fractional integral. His solution,
$f(y)=\frac{d^{\alpha}}{dy^{\alpha}}\mskip 3.0muk(y),$
treats fractional integration and differentiation as inverse operations — a foundational idea of fractional calculus.
Early Fractional Derivative: Lacroix and Leibniz.
The computation of $d^{1/2}(1)/dy^{1/2}$ had precursors. In 1819, Lacroix noted that for $y=x^{m}$,
$\frac{d^{n}}{dx^{n}}x^{m}=\frac{\Gamma(m+1)}{\Gamma(m-n+1)}x^{m-n},\quad n\in\mathbb{N},$
and extended the formula to non-integer $n$. Setting $m=0$, $n=1/2$ gives
$\frac{d^{1/2}}{dx^{1/2}}1=\frac{1}{\sqrt{\pi x}},$
showing that the fractional derivative of a constant is non-zero Lacroix1819 . This result echoes Leibniz’s 1695 speculation that $d^{1/2}x=x\sqrt{dx/x}$ Leibniz1695 — an early hint that differentiation could admit fractional orders.
Fractional Calculus as a Unified Framework.
Abel’s work implicitly defined the Abel integral equation of the first kind:
$_{0}\mathrm{I}_{t}^{\alpha}g(t)=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}\frac{g(\tau)}{(t-\tau)^{1-\alpha}}\mskip 3.0mud\tau=f(t),\quad 0<\alpha<1.$ (15)
Its solution,
$g(t)={}_{0}\mathrm{D}_{t}^{\alpha}f(t),$
assumes the existence of a left–inverse operator satisfying
$_{0}\mathrm{D}_{t}^{\alpha}\mskip 3.0mu{}_{0}\mathrm{I}_{t}^{\alpha}=\mathds{1}.$
This operator — the Riemann–Liouville fractional derivative — was systematized later, but Abel had already used it to solve the tautochrone problem.
Thus, the tautochrone problem served as the first concrete physical motivation for fractional calculus. Abel’s analysis transformed a geometric question (Huygens) into an analytic one, inaugurating a new branch of mathematics that would later find applications across physics, engineering, and applied analysis Mainardi2007 ; Mainardi2007b ; Davis ; Mainardi-Book ; Evangelista2023 ; Oldham1974 ; Podlubny ; Samko ; Yang-Gao-Ju .
IV Fractional Operators: From Riemann–Liouville to Caputo
Fractional calculus extends differentiation and integration to arbitrary (non-integer) orders Lacroix1819 . The emblematic question — what meaning can be given to $d^{1/2}f/dx^{1/2}?$ — was first posed by L’Hôpital to Leibniz in 1695 Leibniz1695 . Today, fractional operators are rigorously defined and widely used in applications.
Riemann–Liouville Operators.
The Riemann–Liouville fractional integral of order $\alpha>0$ is
$_{c}\mathrm{I}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(\alpha)}\int_{c}^{x}\frac{f(t)}{(x-t)^{1-\alpha}}\mskip 3.0mudt.$ (16)
For $\alpha\in(m-1,m)$, the corresponding fractional derivative is defined as
$_{c}\mathrm{D}_{x}^{\alpha}f(x)=\frac{d^{m}}{dx^{m}}\left[{}_{c}\mathrm{I}_{x}^{m-\alpha}f(x)\right],$ (17)
ensuring that $_{c}\mathrm{D}_{x}^{\alpha}\mskip 3.0mu{}_{c}\mathrm{I}_{x}^{\alpha}=\mathds{1}$. These operators satisfy the expected composition laws and reduce to ordinary calculus when $\alpha$ is an integer.
Caputo Operators.
In 1967, Caputo introduced an alternative definition that interchanges the order of differentiation and integration Caputo1967 ; Caputo1969 :
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)={}_{c}\mathrm{I}_{x}^{m-\alpha}\left[\frac{d^{m}}{dx^{m}}f(x)\right],\qquad m-1<\alpha<m.$ (18)
For $0<\alpha<1$, this becomes
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(1-\alpha)}\int_{c}^{x}\frac{f^{\prime}(t)}{(x-t)^{\alpha}}\mskip 3.0mudt.$ (19)
Apart from the normalizing factor $1/\Gamma(1-\alpha)$, Eq. (19) is precisely the operator that appears in Abel’s solution of the tautochrone problem (cf. Eq. (9)). Indeed, Abel’s integral equation
$k=\int_{0}^{y}\frac{f(z)}{(y-z)^{1/2}}\mskip 3.0mudz$
is, up to a constant, a Caputo derivative of order $1/2$.
Relation Between the Two Formulations.
The Riemann–Liouville and Caputo derivatives are connected by
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)={}_{c}\mathrm{D}_{x}^{\alpha}\bigl[f(x)-f(c)\bigr],\qquad 0<\alpha<1.$ (20)
Thus they coincide when $f(c)=0$. For $c\to$-$\infty$, both definitions also coincide provided the function and its derivatives vanish at infinity — a property crucial for modelling stationary processes in fractional dynamical systems, wave propagation, and impedance problems Evangelista2018 ; Mainardi-Book ; Mainardi-C .
Historical Perspective.
While the modern definitions above were systematized in the 20th century, their conceptual roots lie in 19th-century problems. Abel’s 1823–1826 work on the tautochrone implicitly used what we now call the Caputo derivative, and his solution method anticipated the Riemann–Liouville fractional integral. Thus, the tautochrone problem not only motivated the first concrete application of fractional operators but also shaped their early development.
V Concluding Remarks
In this work, we have highlighted a convergence of interests between Huygens and Abel —separated by nearly two centuries, yet united in their engagement with mathematical problems deeply connected to physics. For this analysis, we employed the tautochrone problem — one of the significant isoperimetric problems in the scientific literature — as a guiding thread.
We have outlined how the tautochrone was initially studied by Huygens in the 17th century, motivated by physical considerations related to timekeeping, and how its solution involved the cycloid curve. Later, in the 19th century, Niels Henrik Abel, by considering the inverse problem (reconstructing the curve from the descent time as a function of position), arrived at an integral equation — of the Volterra type — which can be reinterpreted, in more modern terms, as a particular case of a fractional equation.
This transition from a classical physical-geometric problem to an abstract formulation involving generalized integrals is a compelling example of the evolution of mathematical language and the conceptual tools of mathematical physics. It is of particular interest to observe how Abel, already in the early 19th century, worked with integrals of arbitrary order, anticipating aspects of fractional calculus and laying the groundwork for the consolidation of its foundations — whose relevance is attested by the interests of our own time. This relevance lies both in the more formal aspects of mathematical analysis and in the numerous applications of this formalism across various scientific domains.
Acknowledgements.
The Authors are grateful to an anonymous referee for a very careful reading of the manuscript and for the valuable suggestions to improve the whole presentation. L. R. E. gratefully acknowledges the hospitality of the Department of Molecular Sciences and Nanosystems at Ca’ Foscari University of Venice, Italy, and the Department of Materials, Environmental Sciences and Urban Planning (SIMAU) at the Polytechnic University of Marche (UnivPM), Ancona, Italy, where he served as a Visiting Scholar during the preparation of this work.
Appendix A Tangent to a Cycloid
The crucial point here is to consider the tangent BG to the cycloid as parallel to the chord AE of its generating circle, as depicted in Fig. 1, in order to relate the Huygens’s geometric derivation to Abel’s solution to the tautochrone problem. We start by defining
$y=AF$
as the height of point B. Thus, from the right triangle $\Delta DEA$, we may write
$AE=\sqrt{AF\dot{A}D}=\sqrt{2ry}$
and
$\sin\theta=\sin\angle EDA=\frac{2ry}{2r}=\sqrt{\frac{y}{2r}}.$
However, $\theta$ is also the angle the tangent BG makes with the horizontal, which allow us to write
$\sin\theta=\frac{dy}{ds}\quad{\textrm{and}}\quad\frac{dy}{ds}=\sqrt{\frac{y}{2r}}.$
This equation is essentially the same as Eq. 2, thus providing a link between the approaches of Abel and Huygens. In this framework, we can interpret Eq. 2 as stating that a necessary and sufficient condition for the curve $s=f(y)$ to be a tautochrone with quarter-period $T$ is that there exists a circle of radius
$r=\frac{gT^{2}}{\pi^{2}},$
whose tangents are parallel to the corresponding tangents of the tautochrone, having in mind Fig. 1.
Appendix B Why is the Cycloid a Tautochrone?
Let us try to illustrate in simple terms what makes the cycloid a tautochrone in a uniform gravitational field.
We consider a mass released from the rest and try to show that it oscillates in simple harmonic motion along the length of the cycloid. To do this, we recognize that arclength AB in Fig. 1 may be expressed as
$s=4r\sin\theta,$
where, as before, $\theta$ is the angle the tangent to the cycloid at B makes with the horizontal.
Now, the restoring force on a mass sliding along the cycloid (i.e., the component of its weight tangent to the curve), given by
$F=mg\sin\theta=\frac{mg}{4r}s,$
is proportional to its distance (which is measured along the cycloid) from equilibrium. This way, the mass oscillates in simple harmonic motion with a period
$T_{m}=2\pi\sqrt{\frac{4r}{g}}.$
Why this suggests that the cycloid is a tautochrone? The key point is that the arclength BA of the cycloid, as depicted in Fig. 1, is twice the length of the tangent segment BG. Since the latter, BG = $2r\sin\theta$ is proportional to $\sin\theta$, the same occurs with the former, which is precisely the condition for simple harmonic motion.
This indicates that the Abel’s proof says that the only way for a curve to be a tautochrone is that a mass released from the rest oscillates in a simple harmonic motion along the curve.
B.1 The Authors
Luiz Roberto Evangelista
is Professor of Theoretical Physics at the PhD Preogram in Physics at the University of Maringá (UEM), Brazil. His research interest lie in complex fluids, complex systems, and history of physics, and include mathematical physics of liquid-crystals, diffusion problems, and adsorption-desorption phenomena.
Francesco Mainardi
is Professor (Retired) of Mathematical Physics at the University of Bologna, renowned for his pioneering work in fractional calculus. His research focuses on fractional differential equations, viscoelasticity, and anomalous diffusion. He played a key role in promoting the Mittag-Leffler and Wright functions in applied contexts. Mainardi authored influential books, including Fractional Calculus and Waves in Linear Viscoelasticity. He remains an active voice in the field, shaping theory and applications across physics and engineering.
B.2 Financial Support
No specific financial support was required for the work. L. R. E. acknowledges the continuous partial financial support from the Brazilian Agency — CNPq, by means of the Grant n. 30295/2022-9.
B.3 Authors’ Contributions
Both authors contributed equally to the final version of this paper. The initial idea was suggested by F.M. The conceptual framework was developed after a fruitful meeting in Pisa in 2024. The first draft was written by L.R.E. Refinements and the first complete version emerged after an extended and productive exchange of correspondence between the two authors. The current and final form of the paper is the result of a joint effort, consolidated during a personal meeting in Rimini — land of Francesca and Paola Malatesta, whose tragic love was immortalized by Dante in the Divine Comedy — on the shores of the Adriatic Sea.
B.4 Conflict of Interest
The authors declare no potential competing interests.
B.5 Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
B.6 Generative Artificial Intelligence (AI)
This manuscript benefited from the use of Generative AI tools (ChatGPT and DeepSeek) for translation assistance and grammar refinement, given that the authors are not native speakers of English.